OpenStudy (lisa123):

Vertex: (7, −1); point: (5, 11). Find the standard form of the equation

OpenStudy (anonymous):

y= a(x-h)^2+k

OpenStudy (lisa123):

Yes I know the equation I just don't know what to do with the points

OpenStudy (anonymous):

vertex (h,k) Point (x,y)...plug them in

OpenStudy (lisa123):

okay so a(7-1) +7?

OpenStudy (anonymous):

No..look at the values for the vertex and the point and where they are in the equation

OpenStudy (lisa123):

oh a(7-7)^2-1

OpenStudy (anonymous):

Now solve for (a)

OpenStudy (lisa123):

I thought 7 was h and 5 was x how come its not (5+7)^2

OpenStudy (anonymous):

Your right but it is 5-7

OpenStudy (anonymous):

11=a(5-7)^2 -1

OpenStudy (lisa123):

11=a(4) -1 12=4a 3=a

OpenStudy (anonymous):

Good now plug (a) and (h.k) back into the equation y= a(x-h)^2+k

OpenStudy (lisa123):

y= 3(7-5)^2-1 y=12-1 y=11

OpenStudy (anonymous):

just plug in (h,k) and (a)....so you have y= 3(x-7)^2 -1

OpenStudy (anonymous):

so what is (x-7)^2 ?

OpenStudy (lisa123):

oh okay and that's the standard form?

OpenStudy (anonymous):

No that is vertex form....we need to convert to standard form which is y-ax^2+bx+c....so what is (x-7)^2 times 3?

OpenStudy (anonymous):

y=ax^2+bx+c

OpenStudy (lisa123):

y=3x^2+147

OpenStudy (lisa123):

idk how to find b

OpenStudy (anonymous):

(x-7) ^2 is (x-7) (x-7) equals x^2 -14x+49 do you see now multiply by 3

OpenStudy (anonymous):